OpenStudy (anonymous):

Meg wants to take tennis lessons. In the chart, four instructors are listed with what they charge. Which instructor gives the best deal? A) Mark B) Rick C) Andy D) Steve

OpenStudy (mahoganie.carson):

The Best Deal Would Be Rick I Remember My Sister Doing This One:)

OpenStudy (mahoganie.carson):

Here is the problem Mark 1 hour= 100/10 =$10/hr Rick = 80/9 =$8.89/hr Andy =75/8 = $9.38/hr Steve =60/5 = $12/hr